OpenStudy (anonymous):

integration of u^2 sinha(t-u) du

OpenStudy (experimentx):

\[ \int u^2 \sinh a(t-u) \; du\]

OpenStudy (anonymous):

how will i derive

OpenStudy (experimentx):

try integration by parts

OpenStudy (experimentx):

Woops!! this doesn't have closed form in elementry functions http://www.wolframalpha.com/input/?i=integrate+u^2+sinh%28a%28t-u%29%29+du

OpenStudy (anonymous):

but it doesn't show the steps

OpenStudy (anonymous):

and i am little weak in it

OpenStudy (experimentx):

hold on ... i'll check again.

OpenStudy (anonymous):

sure

OpenStudy (experimentx):

you sure that 'a' is there?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

the problem is from the book of laplace transformation by spiegel

OpenStudy (ash2326):

Use Integration by parts \[\int u^2 \sinh (a(t-u)) du\] \[u^2\cosh (a(t-u))-\int \frac{d}{du} u^2 \times \int (\sinh (a(t-u))\] \[u^2\cosh (a(t-u))/(-a)-\int 2x \times \sinh (a(t-u))/(-a)\] Now apply by parts on second term

OpenStudy (experimentx):